OpenStudy (rayford):

solve -5sqrt x-9-7=23

OpenStudy (anonymous):

Is it \[-5\sqrt{x - 9} - 7 = 23\]

OpenStudy (rayford):

yes :)

OpenStudy (anonymous):

ALright. \[-5\sqrt{x - 9} - 7 = 23 \implies -5\sqrt{x - 9} = 30 \implies \sqrt{x - 9} = -6\]At this point, notice how it's negative. Since you have a square root, the solution can never be negative, so it's no solution :)

OpenStudy (rayford):

thank you very much :)

OpenStudy (anonymous):

np :)